# Quantum Corrections to Symmetron Fifth-Force Profiles

### v1

@@ Abstract
Nonlinear scalar-tensor theories of gravity have been considered as candidates for dark matter and dark energy. Often, they possess screening mechanisms which allow the fifth force mediated by the additional scalar degree(s) of freedom to evade detection from local experiments. Their classical behaviour is well studied, but their quantum nature is relatively unexplored. We outline a Green’s function method for obtaining the leading-order quantum corrections to the classical symmetron field profile in the vicinity of a spherically symmetric extended source. For parameters that experiments had previously ruled out, our calculations indicate that the symmetron force may be weaker than the classical field suggests.
@@ 1 Introduction
Scalar-tensor theories of gravity are ubiquitous in modern theoretical physics. Their study is motivated by many of the most popular extensions to the Standard Model — including extra dimensions [1], supersymmetry [2] and string theory [3] — and they make natural candidates to explain observations related to dark matter [4, 5], dark energy [6, 7] and inflation [8]. The scalar fields they introduce mediate new “fifth forces” between matter particles, a phenomenon which cannot be avoided without appealing to scale invariance [9, 10] or conformal invariance [11, 12]. Consequently, measurements of gravity on Earth and in the Solar System would have immediately ruled out theories that give rise to fifth forces were it not for nonlinearities in their field equations, which can screen them from local tests. The precise form of these nonlinearities, and thus screening mechanisms, depends on the specific features of the model. For instance, in theories with screening mechanisms, such as the chameleon model [13], the mass $m$ of the scalar field is background-dependent, becoming high in environments at least as dense as the Solar System and low in environments less dense than interstellar space. Consequently, the range of the fifth force $m^{-1}$ is inversely related to the background density, and local observations remain consistent with the predictions of general relativity. For further details, and a review of other screening mechanisms, see Ref. [14].
@@
The symmetron model [15], which is similar in some ways to the chameleon model, is the scalar field under consideration in this work. In regions of space where the density of matter is higher than some critical value, the vacuum expectation value (VEV) of the field vanishes and the dynamics remain invariant under a $\mathbb{Z}_{2}$ symmetry $\phi\rightarrow-\phi$. In regions of space where the matter density is lower than the critical density, the symmetry is broken, and the field acquires a nonzero VEV. The coupling strength of the symmetron’s fifth force is proportional to its VEV, and so it too is nonzero in low-density regions and negligible in high-density regions. With an appropriate choice of parameters, the symmetron model may take on the characteristics of conventional modified gravity models or even have sufficient energy density to contribute to the gravitational potential of a galaxy as a dark matter component [16]. Searches for symmetrons use a wide variety of techniques, with tests including atom interferometry [17, 18, 19] and measurement of the electron dipole moment [20]. The most recent constraints [21] are derived from experiments using magnetically levitated sensors and improve on previous bounds by over six orders of magnitude. See also Ref. [22] for the recent constraints covering different areas of the parameter space, including forecasts for future experiments.
@@
Despite the obvious upshot of leaving the laws of gravity unchanged exactly where we expect them to be, the nonlinearities in screened gravity models can present significant technical challenges. For instance, unless one makes certain simplifying assumptions, the classical field equations are often very difficult, if not impossible, to solve in closed form. Exact solutions have been found in 1D [23, 24, 25] and are immediately applicable to experiments which effectively vary along one spatial dimension. These include Casimir experiments [26], torsion pendulum experiments [27, 28] and bouncing neutron experiments [29, 30, 31]. Technical difficulty may partially explain why much of the literature deals exclusively with the dynamics of classical fields, as may the conventional wisdom regarding quantum corrections, given formally by Erick Weinberg [32] (see also Ref. [33]). To briefly summarise his argument, one can meaningfully talk about a classical field on some macroscopic length scale $L$ if quantum fluctuations averaged over $L$ are small compared to the variation of the classical field over $L$. Simultaneously, $L$ must be small compared to the characteristic length scale of the classical field configuration. This line of reasoning boils down to an inequality, relating Lagrangian parameters to the length scale $L$, which must be satisfied for the classical field profile not to suffer from large quantum corrections. Applied to “normal” field theories, this argument tends to suggest that quantum corrections become relevant at high energies or, equivalently, small length scales. This reasoning breaks down for very light scalar fields, which have the potential to mediate quantum phenomena over astrophysical distances.
@@
Strong hints towards a need to move beyond tree-level have existed in the fifth force literature for some time. A recent work by Burrage et al. [33] uses Weinberg’s argument and a generalisation of Derrick’s theorem [34] to show that the classical approximation to the symmetron field is insufficient to describe fifth forces due to highly compact sources. An earlier work by Upadhye et al. [35] demonstrates that if one wants the assumption of small quantum corrections to hold for the chameleon, thereby preserving the reliability of predictions at the time, then theoretical considerations place very strong constraints on the mass of the scalar field. More robust treatments of the quantum nature of fifth force models have also emerged in recent years. Some derive quantum forces from path integrals in chameleon and chameleon-like models [36]. Others have used methods developed in open quantum systems to quantise fluctuations on trivial background solutions of the symmetron model [37, 38]. We aim to build upon these results by quantising fluctuations on nontrivial background solutions.
@@
The paper is arranged as follows. A summary of our quantisation method, as well as a Green’s function method for solving the quantum equation of motion, is given in Section 2. It will essentially provide us with a list of things to calculate: the classical field configuration, the coincident Green’s function and the renormalised tadpole contribution. Section 3 contains a calculation of the exact classical field profile around a non-relativistic spherical source of radius $R$, in the asymptotic limit $R\rightarrow\infty$ (analogous to Coleman’s thin-wall approximation for vacuum decay [39, 40]). This limit makes our problem effectively one-dimensional, and so our results are also relevant for experiments with planar geometry, such as CANNEX [26, 41]. Section 4 contains our computation of the Green’s function, in various domains, as well as its coincidence limit. Finally, in Section 5, making use of the Coleman–Weinberg effective potential, we renormalise the tadpole contribution. This step is performed by numerically integrating the frequency-domain Green’s function along with pseudo-counterterms [40]. Once obtained, we compute the leading-order quantum correction, as well as the one-loop field and force profiles.
@@
We find, in agreement with expectations from Ref. [40], that the general feature of quantum corrections is a flattening of the field profile, which leads to a weakening of the force across the parameter space. The strength of this correction grows almost linearly with the strength of the self-coupling. Smaller self-couplings yield one-loop predictions that are mostly in line with tree-level predictions, as expected. However, we note that quantum corrections due to couplings to the Standard Model are expected to become the dominant contribution in this region of parameter space. Furthermore, in contrast to loop corrections on trivial backgrounds, the shift due to loop corrections here depends on position. These are fundamental modifications to the behaviour of the theory and cannot simply be fine-tuned away. The one-loop approximation is technically invalid for non-perturbative couplings. However, the fact that the magnitude of the correction increases with the coupling strength — already to an appreciable degree in the perturbative regime — may change how we interpret current and future constraints on the parameter space. We discuss the implications of these results in more detail in Section 6.
@@
We employ similar techniques previously applied in the context of vacuum decay [40, 42], as well as for the Fubini-Lipatov instanton [43] and Higgs-Yukawa theory [44]. In these contexts, the validity of the quantisation method is contingent on the discrete spectrum of the fluctuation operator being positive definite. For the symmetron, it is reasonable enough to assume that this is true based on the physical properties of the problem. Nevertheless, for completeness, we prove this fact in Appendix A. We opt to compute the Green’s function directly instead of representing it as a sum over eigenfunctions. The process is somewhat involved, so we detail a simpler version of it in Appendix B.
@@ 2 One-Loop Equation of Motion
The starting point for the method that we use is the quantum effective action [45], which will provide a consistent framework within which to determine the leading quantum corrections to the fifth force profiles. In this section, we review the derivation of the quantum effective action, following the approach of Ref. [46], and summarise the resulting one-loop equation of motion needed for our subsequent analysis.
@@
$\left.\frac{\delta\Gamma[\phi]}{\delta J(x)}\right|_{J=\mathcal{J}}=0\>.$ (2.4)
@@
source-dependent in the sense that it is a functional of the external source $\mathcal{J}$ and not the local source $J$. Similarly, the above expression also defines the external source $\mathcal{J}$ as a functional of $\phi$, for which we adopt the notation $\mathcal{J}(x)[\phi]$. Finally, using the expression above, one can show that
@@
To leading order in $\hbar$, the path integral in (2.5) is dominated by the solution which makes the exponent stationary. We denote this solution by $\varphi$, which is defined by
@@
wherein we defined
@@
which is the inverse two-point function, which we will refer to as the fluctuation operator, and where $V$ is the classical potential. The rest of our analysis assumes that the discrete spectrum of this operator is positive definite, a fact that we will prove in Appendix A.
By substituting the expansion above back into the effective action and computing the functional integral over $\tilde{\Phi}(x)$, we obtain
@@
[…]
@@
$\Gamma[\varphi]=S[\varphi]+\hbar\,\Gamma_{1}[\varphi]\>,$ (2.15)
Note that, by following the approach of Ref. [46], the one-point function here is the same as that which appears in the saddle-point approximation to the path integral. In further distinction to the standard method, we need not take the sourceless limit. As a result, the source in constrained such that $\varphi$ solves the quantum equation of motion,
@@
The stationary condition in (2.8) implies that the above may be rendered in the form
@@
At the one-loop level, the quantum equation of motion reads
@@
where the effective potential for the one-loop quantum field is defined by
@@
When expressed in terms of the classical field plus fluctuations, the equation of motion for $\varphi$ reduces to an equation for $\delta\phi$ alone,
@@
which may be readily solved by a convolution
@@
Computing the one-loop correction to the field profile of the symmetron and the resulting one-loop correction to the fifth force is the aim of our work. The integral above gives us a clear order of operations to carry out. First, we calculate the classical field profile $\varphi_{\text{cl}}$. Then, we compute the Green’s function of the fluctuation operator. Finally, we determine the coincidence limit of the Green’s function, regularise it and renormalise it to obtain the tadpole contribution.
@@ 3 Classical Symmetron Model
The first step in our calculation is to compute the static profile of the classical symmetron field in the vicinity of an extended, spherical source. The full equation of motion cannot be solved in closed form, but we can obtain an analytical estimate of the true solution in the limit of large sources.
@@
where $g$ is the determinant of the Einstein frame metric $g_{\mu\nu}$, $R$ is the Ricci scalar, $M_{\text{Pl}}$ is the reduced Planck mass and $\mathcal{L}_{\text{m}}$ denotes the Lagrangian density of matter fields. These fields are minimally coupled to the Jordan-frame metric,
@@
where $\square\equiv\partial^{\mu}\partial_{\mu}$ is the d’Alembertian, $\tilde{T}$ is the trace of the matter energy-momentum tensor in the Jordan frame, and $V_{\text{eff}}$ is the effective potential. Our matter distribution is non-relativistic and pressureless, so we may write $\tilde{T}\approx-\tilde{\rho}$, with $\rho=A(\phi)^{3}\tilde{\rho}$, as its energy density.
The function $V(\phi)$ is chosen so that the effective potential $V_{\text{eff}}$ has a spontaneously broken $\mathbb{Z}_{2}$ symmetry $\phi\rightarrow-\phi$ in the limit $\rho\rightarrow 0$ [47]. It takes the form [33]
@@
This amounts to a tuning of the cosmological constant, which is not relevant to the analysis that follows. The aforementioned characteristics of the effective potential will also constrain the universal coupling $A(\phi)$, for which the choice
$A(\phi)=1+\frac{1}{2M^{2}}\phi^{2}$ (3.6)
is made, where $M$ is a mass scale that controls the strength of the coupling to matter.
@@
See Figure 1 for a plot of this function for both screened and unscreened sources. From it, we can directly observe that for $\rho>\mu^{2}M^{2}$ the field is screened and sits around the minimum at $\phi=0$, whereas for $\rho<\mu^{2}M^{2}$ the field is unscreened and sits in one of the two nonzero minima.
We want to quantise fluctuations about the static classical field configuration, which obeys the equation
@@
We take the extended source to be a sphere of uniform density $\rho_{0}$ and radius $R$, described by the function
@@
making the classical field a positive function, as well as its derivative. The symmetron model also admits solutions which interpolate between the two degenerate minima, leading to the formation of domain walls. Such solutions are of higher energy than the configuration we are solving for, and so will have sub-leading contributions to the quantum corrections. See Refs. [48, 49, 50] for discussion of domain walls in scalar-tensor theories.

### v2

@@ Abstract
Nonlinear scalar-tensor theories of gravity have been considered as candidates for dark matter and dark energy. Often, they possess screening mechanisms, which allow the fifth force mediated by the additional scalar degree(s) of freedom to evade detection from local experiments. Their classical behaviour is well studied, but their quantum nature is relatively unexplored. We outline a Green’s function method for obtaining the leading-order quantum corrections to the classical symmetron field profile, in the vicinity of a spherically symmetric extended source, in the planar limit. For parameters that experiments had previously ruled out, our calculations indicate that the symmetron force may be weaker than the classical field suggests.
@@ 1 Introduction
Scalar-tensor theories of gravity are ubiquitous in modern theoretical physics. Their study is motivated by many of the most popular extensions to the Standard Model — including extra dimensions [51], supersymmetry [33] and string theory [25] — and they make natural candidates to explain observations related to dark matter [10, 54], dark energy [44, 49] and inflation [4, 32]. The scalar fields they introduce mediate new “fifth forces” between matter particles, a phenomenon which cannot be avoided without appealing to scale [61, 34] or conformal invariance [18, 6]. Consequently, measurements of gravity on Earth and in the Solar System would have immediately ruled out theories that give rise to fifth forces were it not for nonlinearities in their field equations, which can screen them from local tests. The precise form of these nonlinearities, and thus screening mechanisms, depends on the specific features of the model. For instance, in theories such as the chameleon model [48], the mass $m$ of the scalar field is background-dependent, becoming high in environments at least as dense as the Solar System and low in environments less dense than interstellar space. Consequently, the range of the fifth force $m^{-1}$ is inversely related to the background density, and local observations remain consistent with the predictions of general relativity. For further details, and a review of other screening mechanisms, see ref. [7].
@@
The symmetron model [42], which is similar in some ways to the chameleon model, is the scalar field under consideration in this work. In regions of space where the density of matter is higher than some critical value, the vacuum expectation value (VEV) of the field vanishes and the dynamics remain invariant under a $\mathbb{Z}_{2}$ symmetry $\phi\rightarrow-\phi$. In regions of space where the matter density is lower than the critical density, the symmetry is broken, and the field acquires a nonzero VEV. The coupling strength of the symmetron’s fifth force is proportional to its VEV, and so it too is nonzero in low-density regions and negligible in high-density regions. With an appropriate choice of parameters, the symmetron model may take on the characteristics of conventional modified gravity models or even have sufficient energy density to contribute to the gravitational potential of a galaxy as a dark matter component [17]. Searches for symmetrons use a wide variety of techniques, with tests including atom interferometry [16, 24, 57] and measurement of the electron magnetic dipole moment [8]. The most recent constraints [67] are derived from experiments using magnetically levitated sensors and improve on previous bounds by over six orders of magnitude. See also ref. [35] for the recent constraints covering different areas of the parameter space, including forecasts for future experiments.
Exact solutions have been found in 1D [15, 56, 21] and are immediately applicable to experiments which effectively vary along one spatial dimension. These include Casimir experiments [60], torsion pendulum experiments [64, 65] and bouncing neutron experiments [29, 14, 13]. The solutions have also found application in the context of the Hubble tension and the cosmic distance ladder. In particular, some authors find that the cosmic distance ladder is a probe comparable in strength to Solar System tests, and that the symmetron generally increases the Hubble tension [43]. Interestingly, if the symmetron does not couple universally to matter and instead has a Yukawa coupling to the electron, its contribution to the electron mass could instead alleviate the tension [62].
@@
Despite the obvious upshot of leaving the laws of gravity unchanged exactly where we expect them to be, the nonlinearities in screened gravity models can present significant technical challenges. For instance, unless one makes certain simplifying assumptions, the classical field equations are often very difficult, if not impossible, to solve in closed form. Technical difficulty may partially explain why much of the literature deals exclusively with the dynamics of classical fields, as may the conventional wisdom regarding quantum corrections, given formally by Erick Weinberg [66] (see also ref. [20]). To briefly summarise his argument, one can meaningfully talk about a classical field on some macroscopic length scale $L$ if quantum fluctuations averaged over $L$ are small compared to the variation of the classical field over $L$. Simultaneously, $L$ must be small compared to the characteristic length scale of the classical field configuration. This line of reasoning boils down to an inequality, relating Lagrangian parameters to the length scale $L$, which must be satisfied for the classical field profile not to suffer from large quantum corrections. Applied to “normal” field theories, this argument tends to suggest that quantum corrections become relevant at high energies or, equivalently, small length scales. This reasoning breaks down for very light scalar fields, which have the potential to mediate quantum phenomena over astrophysical distances.
@@
Strong hints towards a need to move beyond tree-level have existed in the fifth force literature for some time. A recent work by Burrage et al. [20] uses Weinberg’s argument and a generalisation of Derrick’s theorem [30] to show that the classical approximation to the symmetron field is insufficient to describe fifth forces due to highly compact sources. An earlier work by Upadhye et al. [63] demonstrates that if one wants the assumption of small quantum corrections to hold for the chameleon, then theoretical considerations place very strong constraints on the mass of the scalar field. More robust treatments of the quantum nature of fifth force models have also emerged in recent years. Some derive quantum forces from path integrals in chameleon and chameleon-like models [11, 12]. Others have used methods developed in open quantum systems to quantise fluctuations on spatially homogeneous background solutions of the symmetron model [46, 22]. We aim to build upon these results by quantising fluctuations on spatially varying background solutions.
The aim of this paper is to determine an analytical estimate of the relevance of quantum corrections to the symmetron field. This will allow us to make the first estimation of the theoretical uncertainty on fifth-force calculations that arise from neglecting radiative effects. We begin with a spherical problem but, for tractability, we work in its planar limit. This makes the dynamics effectively one-dimensional, and so our results are also relevant for experiments with planar geometry, such as CANNEX [60, 59].
@@
The paper is arranged as follows. A summary of our quantisation method, as well as a Green’s function method for solving the quantum equation of motion, is given in section 2. It will essentially provide us with a list of things to calculate: the classical field configuration, the Green’s function and the renormalised tadpole contribution. Section 3 contains a calculation of the exact classical field profile around a non-relativistic spherical source of radius $R$, in the limit where $R$ is much greater than the field’s Compton wavelength (analogous to Coleman’s thin-wall approximation for vacuum decay [23, 36]). Section 4 contains our computation of the Green’s function, as well as its coincidence limit. This is done in the so-called planar-wall approximation, so that the sum over angular momentum numbers is approximated by an integral over a continuum [36]. Finally, in section 5, making use of the Coleman–Weinberg effective potential, we renormalise the tadpole contribution. This step is performed by numerically integrating the frequency-domain Green’s function along with pseudo-counterterms [36]. Once obtained, we compute the leading-order quantum correction, as well as the one-loop field and force profiles.
@@
We find, in agreement with expectations from ref. [36], that the general feature of quantum corrections is a flattening of the field profile, which leads to a weakening of the force across the parameter space. The strength of this correction grows almost linearly with the strength of the self-coupling. Smaller self-couplings yield one-loop predictions that are mostly in line with tree-level predictions, as expected. However, we note that quantum corrections due to couplings to the Standard Model are expected to become the dominant contribution in this region of parameter space. Furthermore, in contrast to loop corrections on spatially homogeneous backgrounds, the shift due to loop corrections here depends on position. These are fundamental modifications to the behaviour of the theory and cannot simply be fine-tuned away. The one-loop approximation is technically invalid for non-perturbative couplings. However, the fact that the magnitude of the correction increases with the coupling strength — already to an appreciable degree in the perturbative regime — may change how we interpret current and future constraints on the parameter space. We discuss the implications of these results in more detail in section 6.
@@
We employ similar techniques previously applied in the context of vacuum decay [36, 37], as well as for the Fubini-Lipatov instanton [39] and Higgs-Yukawa theory [1]. In these contexts, the fluctuation operator has at most one negative eigenmode, a signal of the expected instability. On the other hand, for static solutions to the symmetron field equations, we expect a positive definite spectrum, and we show this to be the case in appendix A. We opt to compute the Green’s function directly instead of representing it as a sum over eigenfunctions. The process is somewhat involved, so we detail a simpler version of it in appendix B.1 before a more complete analysis in appendix B.2.
@@ 2 One-loop equation of motion
The starting point for the method that we use is the quantum effective action [40], which will provide a consistent framework within which to determine the leading quantum corrections to the fifth-force profiles. In this section, we review the derivation of the quantum effective action, following the approach of ref. [38], and summarise the resulting one-loop equation of motion needed for our subsequent analysis.
@@
$\left.\frac{\delta\Gamma_{J}[\phi]}{\delta J(x)}\right|_{J=\mathcal{J}}=0\>.$ (2.4)
@@
source-dependent in the sense that it is a functional of the external source $\mathcal{J}$. Similarly, the above expression also defines the external source $\mathcal{J}$ as a functional of $\phi$, for which we adopt the notation $\mathcal{J}(x)[\phi]$. Finally, using the expression above, one can show that
@@
To leading order in $\hbar$, the path integral in eq. (2.5) is dominated by the solution that makes the exponent stationary. We denote this solution by $\varphi$, and it is defined by
@@
wherein we defined the inverse two-point function
@@
which we will refer to as the fluctuation operator, where $V$ is the classical potential and $\square=\partial_{\mu}\partial^{\mu}$ is the d’Alembertian. The rest of our analysis assumes that the discrete spectrum of this operator is positive definite, which we will prove in appendix A.
By substituting the expansion in eq. (2) back into the effective action and computing the functional integral over $\tilde{\Phi}(x)$, we obtain
@@
We now eliminate $\phi$ in favour of $\varphi$ by writing $\varphi(x)=\phi(x)-\hbar\,\delta\varphi(x)$ and expanding the left-hand side as
@@
$\Gamma[\varphi]=S[\varphi]+\hbar\,\Gamma_{1}[\varphi]\>.$ (2.15)
Note that, by following the approach of ref. [38], the one-point function here is the same as that which appears in the saddle-point approximation to the path integral. As a result, the source is constrained such that $\varphi$ solves the quantum equation of motion,
@@
The stationarity condition in eq. (2.8) implies that the above may be rendered in the form
@@
At the one-loop level, the quantum equation of motion therefore reads
@@
where the derivative effective potential for the one-loop quantum field is defined by
@@
When expressed in terms of the classical field to first order in $\delta\phi$, the equation of motion for $\varphi$ reduces to an equation for $\delta\phi$ alone,
@@
which may readily be solved as
@@
Computing the one-loop correction to the field profile of the symmetron and the resulting one-loop correction to the fifth force is the aim of this work. The integral above gives us a clear order of operations to carry out. First, we calculate the classical field profile $\varphi_{\text{cl}}$. Then, we compute the Green’s function of the fluctuation operator. Finally, we determine the coincidence limit of the Green’s function, regularise it and renormalise it to obtain the tadpole contribution, and use eq. (2.23) to determine the leading quantum correction to the classical field profile.
@@ 3 Classical symmetron model
The first step in the calculation is to compute the static profile of the classical symmetron field in the vicinity of an extended, spherical source. The full equation of motion cannot be solved in closed form, but we can obtain an analytical estimate of the true solution in the limit of large sources.
3.1 Theoretical background
@@
where $g$ is the determinant of the Einstein-frame metric $g_{\mu\nu}$, $R$ is the Ricci scalar, $M_{\text{Pl}}$ is the reduced Planck mass and $\mathcal{L}_{\text{m}}$ denotes the Lagrangian density of matter fields. These fields are minimally coupled to the Jordan-frame metric,
@@
$\tilde{T}$ is the trace of the matter energy-momentum tensor in the Jordan frame, and $V_{\text{eff}}$ is the effective potential. Our matter distribution is non-relativistic and pressureless, so we may write $\tilde{T}\approx-\tilde{\rho}$, with $\rho=A(\phi)^{3}\tilde{\rho}$, as its energy density.
The functions $V(\phi)$ and $A(\phi)$ are chosen so that the effective potential $V_{\text{eff}}$ has a spontaneously broken $\mathbb{Z}_{2}$ symmetry $\phi\rightarrow-\phi$ in the limit $\rho\rightarrow 0$ [41]. The potential $V(\phi)$ takes the form [20]
@@
This amounts to a tuning of the cosmological constant, which is not relevant to the analysis that follows.
For the universal coupling $A(\phi)$ we have
$A(\phi)=1+\frac{1}{2M^{2}}\phi^{2}\>,$ (3.6)
where $M$ is a mass scale that controls the strength of the coupling to matter.
@@
See figure 1 for a plot of this function for both screened and unscreened sources. From the figure, we can directly observe that for $\rho>\mu^{2}M^{2}$ the field is screened and sits around the minimum at $\phi=0$, whereas for $\rho<\mu^{2}M^{2}$ the field is unscreened and sits in one of the two nonzero minima.
We want to solve for the static classical field configuration, which obeys the equation
@@
We take the extended matter source to be a sphere of uniform density $\rho_{0}$ and radius $R$, described by the function
@@
making the classical field a positive function, as well as its derivative. The symmetron model also admits solutions which interpolate between the two degenerate minima, leading to the formation of domain walls. Such solutions are of higher energy than the configuration that we are solving for, and so will have sub-leading contributions to the quantum corrections. See refs. [55, 50, 26] for discussions of domain walls in scalar-tensor theories.